How much of an alloy that is 10% copper should be mixed with 300 ounces of an
alloy that is 90% copper in order to get an alloy that is 50% copper?




___ ounces

Answer by checkley71(8403)

.1X+.90*300=.5(300+X)
.1X+270=150+.5X
.1X-.5X=150-270
-.4X=-120
X=-120/-.4
X=300 OUNCES OF 10% COPPER IS REQUIRED.
PROOF:
.10*300+.9*300=.5*600
30+270=300
300=300